A courier company has two vehicles available for local deliveries. Because of the demand on their time and 

the chance of mechanical failure, the probability that a specific vehicle will be available when needed is 

0.85. The availability of one vehicle is independent of the other.

4.1 In the event of large orders, what is the probability that both vehicles will be available?

4.2 What is the probability that none of the vehicles will be available?


1
Expert's answer
2021-11-08T19:34:36-0500

1.

"P=0.85\cdot 0.85=0.7225"


2.

"P=(1-0.85)(1-0.85)=0.0225"